Using Equations to solve world problems.. Please explain how to solve and the answer..
1. George had 9 times as many stamps in his collection as Daniel has. Bryan has 2 times as many stamps as Daniel. The combined stamp collection of the three boys is 144. How many stamps does each boy have?

I need to Write the equation:
Solve for "n" :
Answer:
Thanks

Answer by checkley71(8403)

G=9D B=2D AND D=D G+B+D=144 THEN 9D+2D+D=144 OR 12D=144 THEN D=144/12 OR D=12
DANIEL HAS 12 STAMPS, GEORGE HAS 9*12 OR 108 STAMPS AND BRYAN HAS 2*12 OR 24
STAMPS THUS THE TOTAL STAMPS IS 12+108+24 OR 144 STAMPS.